# Scaling Optimal LR Across Token Horizons

### v2

@@ Abstract
State-of-the-art LLMs are powered by scaling – scaling model size, dataset size, and cluster size. It is economically infeasible to extensively tune hyperparameters for the largest runs. Instead, approximately optimal hyperparameters must be inferred or transferred from smaller experiments. Hyperparameter transfer across model sizes has been studied in Yang et al. (2022). However, hyperparameter transfer across dataset size – or token horizon – has not been studied yet. To remedy this we conduct a large-scale empirical study on how optimal learning rate (LR) depends on the token horizon in LLM training. We first demonstrate that the optimal LR changes significantly with token horizon – longer training necessitates smaller LR. Secondly, we demonstrate that the optimal LR follows a scaling law and that the optimal LR for longer horizons can be accurately estimated from shorter horizons via such scaling laws. We also provide a rule-of-thumb for transferring LR across token horizons with zero overhead over current practices. Lastly, we provide evidence that LLama-1 used too high LR, and argue that hyperparameter transfer across data size is an overlooked component of LLM training.
@@ 1 Introduction
State-of-the-art LLMs are scaled in multiple dimensions. The models are becoming increasingly large, e.g. Grok-1.5 has 314 billion (B) parameters (xAI, 2024). The clusters used to train them are growing in size, e.g. the recently operational Memphis super-cluster contains over 100,000 H100 GPUs (Alcorn, 2024). Lastly, the training datasets are growing, e.g. LLama-3 was trained on 15 Trillion (T) tokens (Dubey et al., 2024). At these scales, it is infeasible to extensively tune hyperparameters. Practitioners must instead resort to hyperparameter transfer, a process where approximately optimal hyperparameters for large-scale experiments are inferred from experiments at a smaller scale. Perhaps the most famous work on hyperparameter transfer is muP (Yang et al., 2022) – a methodology for transferring optimal hyperparameter from a small model to a large model. While hyperparameter transfer across model size is a well-studied problem, transfer across dataset size remains understudied. This paper aims to remedy this shortcoming in the literature.
In this paper, we present a large-scale study on hyperparameter transfer across dataset sizes. We specifically focus on learning rate (LR), an important hyperparameter that influences training stability, generalization, and convergence speed. Our study is essentially a large ablation experiment where we vary LR and token horizon for a few different LLM models. We consider $>$250 training runs in total. To keep the scope and computational requirements manageable we focus on how LR depends on the token horizon, and only present some preliminary results on its interaction on model size. Our experiments are based on standard public resources – the Megatron codebase (Shoeybi et al., 2019) and the RefinedWeb dataset (Penedo et al., 2023). Our three main contributions are:
@@ 3 Experiments
Experimental Setup. Our experimental setup closely follows the setup of the GPT-3 paper (Brown, 2020). We consider model sizes of 50 million (m), 125m, 350m, 760m, 1.3 billion (B), and 2.7B parameters using the architectures of Table 2.1 in the GPT-3 paper. The table is replicated as Table 4 in Appendix A and also lists the LRs used in the GPT-3 paper. We use hyperparameters following GPT-3 – weight decay of 0.1, gradient clipping of 1.0, and cosine learning decay schedule. The full list of hyperparameters can be viewed in Table 3 in Appendix A. The LR will vary with the training step since we use an LR schedule, and we will use the term LR to refer to the largest LR of a run. We make three adjustments compared to the GPT-3, mainly to prevent divergence of the model for high LRs. 1) For warmup we use the maxima of 1% of the training steps following Muennighoff et al. (2024) and 1000. Having too short a warmup stage is known to cause divergence (Wortsman et al., 2023) which we want to avoid for short training runs. 2) We use qk-norm following Dehghani et al. (2023), this is known to prevent divergence without hurting validation loss (Wortsman et al., 2023). 3) The original GPT-3 paper uses different batch sizes for different model sizes. To avoid confounders we simply use the same batch size of 0.5m tokens all model. We use the RefinedWeb dataset (Penedo et al., 2023), a common-crawl derived dataset of roughly 600B tokens which is known to be of high quality (Penedo et al., 2024). Experiments are run on the Megatron codebase (Shoeybi et al., 2019). Unless mentioned we will use the same seed for all runs. For curve fitting we use least-square fitting with either a first or second-degree polynomial using Numpy and Scipy (Harris et al., 2020).
@@ 3.1 Ablations
In our first experiment, we consider the 350m LLM model from Table 4 in Appendix A. We perform an ablation study where we vary the LR and token horizon and measure the final validation loss. We consider $\{25,50,100,200,400\}$ billion tokens. We will start with the LR of $3\!\times\!10^{-4}$ from Table 4 which is what the GPT-3 paper used (Brown, 2020). We then multiply this base LR by factors $\{0.25,0.5,1,2,4\}$. The LRs we consider are thus $\{7.5\!\times\!10^{-5},1.5\!\times\!10^{-4},3\!\times\!10^{-4},6\!\times\!10^{-4},1.2\!\times\!10^{-3}\}$. For each combination of LR and token horizon, we train a model and record its final validation loss. To make sure we have the best possible resolution around the minima we find the optimal learning rate $LR^{*}$, and then further train with the LRs halfway between $LR^{*}$ and its two nearest neighbors – a procedure we repeat twice. From these losses, we fit a curve that estimates how the final validation loss depends on the LR. For each token horizon we fit a second degree polynomial in $\log(LR)$ which provides an excellent fit. The $R^{2}$ of the fit are 0.995 or better, see Table 6 in the Appendix for exact values. The validation losses and fitted curve are shown in Figure 2. We also estimate the optimal LR by taking the minimizer of the fitted curve. From Figure 2 we make two observations: 1) the optimal LR decreases with longer token horizons, and 2) the quadratic fit provides excellent agreement with experimental data.
@@ 3.2 Scaling Laws
Here $B$ and $\beta$ are two constants that might e.g. depend on the model architecture. Taking the logarithm of both sides of Equation 1 we get
@@
We thus have a linear equation in the unknowns $\log B,\beta$, and can fit these to our experimental results with least squares. Specifically, we use the minimizer of the quadratic fits of Figure 3 as the data points $LR^{*}$. Fitting Equation 2 gives good fits – the curves for four models are shown in Figure 4. The $R^{2}$ of these fits are in the range 0.99-0.96 (see Table 5 in Appendix A for exact values), a rather high number. In Figure 11 in the Appendix we also show the fits for the 1.3B and 2.7B models, which also show a good fit.
To evaluate the scaling laws we cannot solely rely on $R^{2}$ as that is essentially a measure of training loss. We instead need to evaluate the fit on some held-out data. To simulate hyperparameter transfer we thus consider fitting the constants $\log B,\beta$ on token horizons $25B,50B$ and $100B$ – and then use these constants to predict the optimal LR at longer horizons. We will use the optimal LR we have empirically found with quadratic fits as the correct LR. The results are illustrated for the 50m model in Table 1. We see a reasonably good fit with a relative error of 10-15% – a clear improvement over not scaling LR at all which has a relative error of $>250\%$. We thus conclude that it is possible to perform hyperparameter transfer across token horizons with our scaling laws. Table 1 is repeated for more models with similar fits in Appendix B.
@@
Table 1: Simulated hyperparameter transfer across token horizons for a 50m model. We use the empirically measured optimal LR at 25,50 and 100B tokens to estimate the optimal LR at longer horizons by fitting the constants of Equation 2. We find a reasonable fit when scaling up the horizon to 800B tokens, with a relative error of 10-15 %. The relative error of using the best LR at the 100B horizon for an 800B horizon is $>$ 250%.
@@ 4 Scaling Law with respect to model size
Figure 7: Fit of Equation 3 compared to the experimental results. The data points for the 7B model of Section 4.1 are excluded at the time of fitting and used as validation data. We have an $R^{2}$ of $0.978$ on this validation data.
@@ 4.1 A Case-study on Llama-1
Figure 8: The optimal LR of the LLama-1 7B model as a function of dataset size. We find that $1.15\!\times\!10^{-4}$ is optimal for 1T tokens whereas the Llama-1 paper used $3\!\times\!10^{-4}$.
@@ 5 Related Work
LR has long been known as an important hyperparameter, and hence there is ample work on how to select it (Goyal, 2017). Larger models are known to require smaller LR, and e.g. Kaplan et al. (2020) suggests the formula $LR(N)\approx 0.003239-0.0001395\log(N)$ for tuning LR as a function of model size. MuP (Yang et al., 2022) is a principled methodology for selecting LR when the model is scaled, and the method is actively researched (Lingle, 2024; Everett et al., 2024; Noci et al., 2024; Blake et al., 2024). Recently Everett et al. (2024) showed that LR transfer across model size is possible both with different parametrizations and optimizers. However, a limitation in both Yang et al. (2022) and Everett et al. (2024) is the fixed training horizon assumption used in the theoretical derivations. Our work directly explores this limitation. One notable conclusion in Everett et al. (2024) is that extrapolating a LR to larger models may significantly overestimate optimal LR in the compute optimal setting. We reach a different conclusion, the exponent for model size is independent of the token horizon and vice versa. While their results might hold true for other optimizers, a closer look at Table 11 in Everett et al. (2024) shows that for the Adam optimizer, the scaling exponents are similar for the two token horizons they considered. At last, we mention that Bi et al. (2024) recently fits the optimal LR as a function of total compute (which will combine batch size, model size, and total training duration) for two small models.
@@ 6 Discussion
Limitations. To limit the scope and computational requirements of our study we have intentionally focused on a narrow area – scaling token horizons and changing LR with otherwise fixed LLM recipes. With this limited scope, there are naturally many limitations to our study. We have only extended the scaling laws to roughly 800B tokens, while many SOTA LLMs are trained significantly longer (Dubey et al., 2024). It is well-known that optimal LR depends on model size, and this study has only scratched the surface of this important topic. Beyond just model size, model architectural modifications like mixture of experts (Shazeer et al., 2017), different attention types (Ainslie et al., 2023), and state-space models (Gu & Dao, 2023) could plausibly interact with both the LR and token horizon. Other dimensions to consider are additional hyper parameters (Bjorck et al., 2021) and multi-modality (Huang et al., 2023). We leave these topics for future work.

### v3

@@ Abstract
State-of-the-art LLMs are powered by scaling – scaling model size, training tokens, and cluster size. It is economically infeasible to extensively tune hyperparameters for the largest runs. Instead, approximately optimal hyperparameters must be inferred or transferred from smaller experiments. Hyperparameter transfer across model sizes has been studied in Yang et al. (2022). However, hyperparameter transfer across training tokens – or token horizon – has not been studied yet. To remedy this we conduct a large-scale empirical study on how optimal learning rate (LR) depends on the token horizon in LLM training. We first demonstrate that the optimal LR changes significantly with token horizon – longer training necessitates smaller LR. Secondly, we demonstrate that the optimal LR follows a scaling law and that the optimal LR for longer horizons can be accurately estimated from shorter horizons via such scaling laws. We also provide a rule-of-thumb for transferring LR across token horizons with zero overhead over current practices. Lastly, we provide evidence that LLama-1 used too high LR, and thus argue that hyperparameter transfer across data size is an overlooked component of LLM training.
@@ 1 Introduction
State-of-the-art LLMs are scaled in multiple dimensions. The models are becoming increasingly large, e.g. Grok-1.5 has 314 billion (B) parameters (xAI, 2024). The clusters used to train them are growing in size, e.g. the recently operational Memphis super-cluster contains over 100,000 H100 GPUs (Alcorn, 2024). Lastly, the training datasets are growing, e.g. LLama-3 was trained on 15 Trillion (T) tokens (Dubey et al., 2024). At these scales, it is infeasible to extensively tune hyperparameters. Practitioners must instead resort to hyperparameter transfer, a process where approximately optimal hyperparameters for large-scale experiments are inferred from experiments at a smaller scale. Perhaps the most famous work on hyperparameter transfer is muP (Yang et al., 2022) – a methodology for transferring optimal hyperparameter from a small model to a large model. While hyperparameter transfer across model size is a well-studied problem, transfer across token horizon remains understudied. This paper aims to remedy this shortcoming in the literature.
In this paper, we present a large-scale study on hyperparameter transfer across token horizons, we define the latter as the total number of tokens seen during training. We specifically focus on learning rate (LR), an important hyperparameter that influences training stability, generalization, and convergence speed. Our study is essentially a large ablation experiment where we vary LR and token horizon for a few different LLM models. We consider $>$250 training runs in total. To keep the scope and computational requirements manageable we focus on how LR depends on the token horizon, and only present some preliminary results on its interaction on model size. Our experiments are based on standard public resources – the Megatron codebase (Shoeybi et al., 2019) and the RefinedWeb dataset (Penedo et al., 2023). Our three main contributions are:
@@ 3 Experiments
Experimental Setup. Our experimental setup closely follows the setup of the GPT-3 paper (Brown, 2020). We consider model sizes of 50 million (m), 125m, 350m, 760m, 1.3 billion (B), and 2.7B parameters using the architectures of Table 2.1 in the GPT-3 paper. The table is replicated as Table 4 in Appendix A and also lists the LRs used in the GPT-3 paper. We use hyperparameters following GPT-3 – weight decay of 0.1, gradient clipping of 1.0, and cosine learning decay schedule. The full list of hyperparameters can be viewed in Table 3 in Appendix A. The LR will vary with the training step since we use an LR schedule, and we will use the term LR to refer to the largest LR of a run. We make three adjustments compared to the GPT-3, mainly to prevent divergence of the model for high LRs. 1) For warmup we use the maxima of 1% of the training steps following Muennighoff et al. (2024) and 1000. Having too short a warmup stage is known to cause divergence (Wortsman et al., 2023) which we want to avoid for short training runs. 2) We use qk-norm following Dehghani et al. (2023), this is known to prevent divergence without hurting validation loss (Wortsman et al., 2023). 3) The original GPT-3 paper uses different batch sizes for different model sizes. To avoid confounders we simply use the same batch size of 0.5m tokens all model. We use the RefinedWeb dataset (Penedo et al., 2023), a common-crawl derived dataset of roughly 600B tokens which is known to be of high quality (Penedo et al., 2024). Experiments are run on the Megatron codebase (Shoeybi et al., 2019). Unless mentioned we will use the same seed for all runs. For curve fitting we use least-square fitting with either a first or second-degree polynomial using Numpy and Scipy (Harris et al., 2020), and we always have more data points than free parameters in the fits.
@@ 3.1 Ablations
In our first experiment, we consider the 350m LLM model from Table 4 in Appendix A. We perform an ablation study where we vary the LR and token horizon and measure the final validation loss. We consider $\{25,50,100,200,400\}$ billion tokens. We will start with the LR of $3\!\times\!10^{-4}$ from Table 4 which is what the GPT-3 paper used (Brown, 2020). We then multiply this base LR by factors $\{0.25,0.5,1,2,4\}$. The LRs we consider are thus $\{7.5\!\times\!10^{-5},1.5\!\times\!10^{-4},3\!\times\!10^{-4},6\!\times\!10^{-4},1.2\!\times\!10^{-3}\}$. For each combination of LR and token horizon, we train a model and record its final validation loss. To make sure we have the best possible resolution around the minima we find the optimal learning rate $LR^{*}$, and then further train with the LRs halfway between $LR^{*}$ and its two nearest neighbors – a procedure we repeat twice. From these losses, we fit a curve that estimates how the final validation loss depends on the LR. For each token horizon we fit a second-degree polynomial in $\log(LR)$, using a quadratic polynomial as it provides an excellent fit and is the simplest polynomial with a well-defined minimum. The $R^{2}$ of the fit are 0.995 or better, see Table 6 in the Appendix for exact values. The validation losses and fitted curve are shown in Figure 2. We also estimate the optimal LR by taking the minimizer of the fitted curve. From Figure 2 we make two observations: 1) the optimal LR decreases with longer token horizons, and 2) the quadratic fit provides excellent agreement with experimental data.
@@ 3.2 Scaling Laws
Here $B$ and $\beta$ are two constants independent of $D$ that might e.g. depend on the model architecture. Taking the logarithm of both sides of Equation 1 we get
@@
We thus have a linear equation in the unknowns $\log B,\beta$, and can fit these to our experimental results with least squares. Specifically, we use the minimizer of the quadratic fits of Figure 3 as the data points $LR^{*}$. Fitting Equation 2 gives good fits – the curves for four models are shown in Figure 4. The $R^{2}$ of these fits are in the range 0.99-0.96 (see Table 5 in Appendix A for exact values), a rather high number. In Figure 13 in the Appendix we also show the fits for the 1.3B and 2.7B models, which also show a good fit. In Figure 11 in the Appendix we repeat these experiments on a small scale by using the Llama architecture on the smallest 50m model, and find that the $\beta$ transfer well across architectures when using the same model size. In Figure 12 we repeat these experiments by doing multiple epochs over a fixed dataset of 25B tokens, and observe similar scaling behavior.
To evaluate the predictive power of the scaling laws we cannot solely rely on $R^{2}$ as that is essentially a measure of training loss. We instead need to evaluate the fit on some held-out data. To simulate hyperparameter transfer we thus consider fitting the constants $\log B,\beta$ on token horizons $25B,50B$ and $100B$ – and then use these constants to predict the optimal LR at longer horizons. We will use the optimal LR we have empirically found with quadratic fits as the correct LR. The results are illustrated for the 50m model in Table 1. We see a reasonably good fit with a relative error of 10-15% – a clear improvement over not scaling LR at all which has a relative error of $>250\%$. We thus conclude that it is possible to perform hyperparameter transfer across token horizons with our scaling laws. Table 1 is repeated for more models with similar fits in Appendix B.
@@
Table 1: Simulated hyperparameter transfer across token horizons for a 50m model. We use the empirically measured optimal LR at 25,50 and 100B tokens to estimate the optimal LR at longer horizons by fitting the constants of Equation 2. We find a reasonable fit when scaling up the horizon to 800B tokens, with a relative error of 10-15 %. The relative error of using the best LR at the 100B horizon for an 800B horizon is $>$ 250%. The scaling laws thus have predictive power.
@@ 4 Scaling Law with respect to model size
Figure 7: Fit of Equation 3 compared to the experimental results. The data points for the 7B model of Section 4.1 are excluded at the time of fitting and used as validation data. We have an $R^{2}$ of $0.978$ on this validation data. Note, the 7B model uses the Llama architecture while the other data points use the GPT-3 architecture. This experiment thus demonstrates that the scaling laws have predictive power across model architectures.
@@ 4.1 A Case-study on Llama-1
Figure 8: The optimal LR of the LLama-1 7B model as a function of token horizon $D$. We find that $1.15\!\times\!10^{-4}$ is optimal for 1T tokens whereas the Llama-1 paper used $3\!\times\!10^{-4}$.
@@ 5 Related Work
LR has long been known as an important hyperparameter, and hence there is ample work on how to select it (Goyal, 2017). Larger models are known to require smaller LR, and e.g. Kaplan et al. (2020) suggests the formula $LR(N)\approx 0.003239-0.0001395\log(N)$ for tuning LR as a function of model size. MuP (Yang et al., 2022) is a principled methodology for selecting LR when the model is scaled, and the method is actively researched (Lingle, 2024; Everett et al., 2024; Noci et al., 2024; Blake et al., 2024). Recently Everett et al. (2024) showed that LR transfer across model size is possible both with different parametrizations and optimizers. However, a limitation in both Yang et al. (2022) and Everett et al. (2024) is the fixed training horizon assumption used in the theoretical derivations. Our work directly explores this limitation. One notable conclusion in Everett et al. (2024) is that extrapolating a LR to larger models may significantly overestimate optimal LR in the compute optimal setting. We reach a different conclusion, the exponent for model size is independent of the token horizon and vice versa. While their results might hold true for other optimizers, a closer look at Table 11 in Everett et al. (2024) shows that for the Adam optimizer, the scaling exponents are similar for the two token horizons they considered. Another related paper is Wang & Aitchison (2024), which studies how to set the weight decay of Adam when model and dataset is scaled. At last, we mention that Bi et al. (2024) recently fits the optimal LR as a function of total compute (which will combine batch size, model size, and total training duration) for two small models.
@@ 6 Discussion
Limitations. To limit the scope and computational requirements of our study we have intentionally focused on a narrow area – scaling token horizons and changing LR with otherwise fixed LLM recipes. With this limited scope, there are naturally many limitations to our study. We have only extended the scaling laws to roughly 800B tokens, while many SOTA LLMs are trained significantly longer (Dubey et al., 2024). It is well-known that optimal LR depends on model size, and this study has only scratched the surface of this important topic. Beyond just model size, model architectural modifications like mixture of experts (Shazeer et al., 2017), model width and depth, different attention types (Ainslie et al., 2023), and state-space models (Gu & Dao, 2023) could plausibly interact with both the LR and token horizon. The number of repeated tokens can also play a role, and we only provide small-scale experiments on these theme in Figure 12 in the Appendix. Other dimensions to consider are additional hyperparameters such as weight decay which interacts with LR (Bjorck et al., 2021), LR schedules, and multi-modality (Huang et al., 2023). For computational reasons, we defer these topics to future work.
@@ Appendix B Additional Experimental Data
Figure 11: Experiments on the 50m model using the Llama-1 architecture (RMSnorm, Rope embeddings, and so on) of Touvron et al. (2023). Similar to the GPT-3 architecture, we see that optimal LR decreases with the token horizon. On the right we plot the scaling law, using the same $\beta$ as the GPT-3 model. We see an excellent fit, demonstrating that the $\beta$ transfers well across model architectures.
Figure 12: We sample a set of 25B unique tokens, and ablate LR when the training horizon is 25B, 50B and 100B tokens. This corresponds to one, two, or four epochs over the data. We use the 50m model. (Left) The optimal LR decreases when we increase the total token horizon, thus the total number of unique tokens seen does not solely determine the optimal LR. (Right) The slope between the optima in this setting fits the scaling curve from training only only unique tokens. This suggests that it is the token horizon rather than the number of unique tokens which determines the scaling.